OpenStudy (frostbite):

How can we prove the triple product rule from the implicit function theorem, two variables?

OpenStudy (frostbite):

Triple product rule: \[\frac{ \partial x }{ \partial y }\frac{ \partial y }{ \partial z }\frac{ \partial z }{ \partial x }=-1\] implicit function theorem, two variables: see attachment

OpenStudy (frostbite):

@amistre64

OpenStudy (amistre64):

does triple product refer to scalar or vector \[scalar:a\cdot(b\times c)\]\[vector:a\times(b\times c)\]

OpenStudy (amistre64):

i assume since its 1 that its scalar

OpenStudy (frostbite):

Well perhaps, I have no idea where to start, I was just told to give it a shot, but no idea where to start. And i sepose you mean -1?

OpenStudy (amistre64):

yeah, -1 ;) http://en.wikipedia.org/wiki/Triple_product_rule i get this as a start from the wiki, and i am reading up on some other finds; but at the moment i got no idea :)

OpenStudy (frostbite):

Welll i can try uploade the full pdf I have. Don't think it is going to be any better however.

OpenStudy (frostbite):

p. 9 - 17.

OpenStudy (amistre64):

yeah, i cant make much sense of it; either with the google or with the pdf :/

OpenStudy (amistre64):

are we able to rewrite the partials in terms of df/dx, df/dy, df/dz ? cause i tend to see alot of those being tossed about in there

OpenStudy (amistre64):

and it might have something to do with a Jacobian ....

OpenStudy (frostbite):

As far I have been told we may use any known differential notation and calculus.

OpenStudy (amistre64):

http://answers.yahoo.com/question/index?qid=20090319011005AAa0Djq can you make this link out? it seems to address alot of the stuff

OpenStudy (amistre64):

(dF/dx) / (dF/dy) dF dy -- -- = dy/dx dx dF

OpenStudy (amistre64):

dropped a negative again

OpenStudy (frostbite):

Well i don't see how he/she get the "dy/dx = -(∂F/∂x)/(∂F/∂y)....." part is that a calculus rule?

OpenStudy (amistre64):

i think thats the implicit function thrm, or at least a consequence of it

OpenStudy (frostbite):

Yea was just thinking the same. and then you do that for all veriables and just multiply them?

OpenStudy (amistre64):

yep, thats what it looks like to me; granted i cant verify it, but it reads good to me

OpenStudy (frostbite):

Hmm if it looks fine to you i trust your jodgement. Thanks a lot, seems alot more easy, just wondring why they did not write that consequence in the pdf.

OpenStudy (amistre64):

http://www.youtube.com/watch?v=xtgTckGMuWE oh, and patrick has the best explanations